Kirsty manages a restaurant. Each day she buys watermelons and mangoes in the ratio of 3∶2. How many watermelons did she buy if, on one day, the total number of watermelons and mangoes was 200?


1
Expert's answer
2021-10-25T17:59:50-0400

watermelon to mangoes= 3: 2

3+2=5

number of watermelon =( 3/5 ) x 200 = 120